Exercises 263 ted below is information for Zhou Co. for the month of January 2014 E5-6 $212,000 Rent expense Cost of goods sold Freight-out Insurance expense Prepare an income statement s 32.000 d cokulate profabliny 7,000 Sales discounts 12,000 Sales returns and allowances 20,000 370,000 wages expense 60,000 Sales revenue an income statement using the format presented on page 245, Assume a 25% e the profit margin and the gross profit rate. tax rate. sFinancial information is presented here for two companies Compute missing amounts and calculate profitability ratios. Yanik $90,000 84,000 Nunez Company Company Sales revenue Sales returns and allowances Net sales Cost of goods sold Gross profit Operating expenses Net income S 5,000 100,000 40,000 17,000 58,000 14,380 Instructions (a) Fill in the missing amounts. Show all computations. b) Calculate the profit margin and the gross profit rate for each company. (c) Discuss your findings in part (b). E5-8 In its income statement for the year ended December 31, 2014, Gavin Company Prepare multiple-step income reported the following condensed data. statement and calculate profitability ratios Salaries and wages Loss on disposal of plant $465,000 assets 987,000 Sales revenue $ 83,500 (04,6, AP 2,210,000C expense Cost of goods sold Interest expense Interest revenue Depreciation expense 25,000 XLS 160,000 110,000 71,000 65,000 Income tax expense Sales discounts 310,000 Utilities expense
ancial information is presented here for two companies. E5-7 Compute missing amo and calculate profitab ratios. Yanik Company $90,000 Nunez Company L0 4,6), AP Sales revenue Sales returns and allowances Net sales Cost of goods sold Gross profit Operating expenses Net income 5,000 84,000 58,000 100,000 40,000 17,000 14,380 Instructions (a) Fill in the missing amounts. Show all computations. (b) Calculate the profit margin and the gross profit rate for each company (c) Discuss your findings in part (b).
Answer #1

Answer

  • All working forms part of the answer
  • You have asked for E5 – 7
  • Requirement ‘a’

Yanik Company

Nunez Compnay

A

Sales revenue

$         90,000.00

$       105,000.00

B

Sales return & Allowances

$            6,000.00

$           5,000.00

C = A - B

Net Sales

$         84,000.00

$       100,000.00

D

Cost of Goods Sold

$         58,000.00

$         60,000.00

E = C- D

Gross Profit

$         26,000.00

$         40,000.00

F

Operating Expenses

$         14,380.00

$         23,000.00

G = E - F

Net Income

$         11,620.00

$         17,000.00

--Working

Yanik Company

Nunez Compnay

A

Sales revenue

90000

=100000+5000

B

Sales rerurn & Allowances

=90000-84000

5000

C = A - B

Net Sales

84000

100000

D

Cost of Goods Sold

58000

=100000-40000

E = C- D

Gross Profit

=84000-58000

40000

F

Operating Expenses

14380

=40000-17000

G = E - F

Net Income

=26000-14380

17000

  • Requirement ‘b’

Working

Yanik Company

Nunez Compnay

A

Net Income

$         11,620.00

$         17,000.00

B

Net Sales

$         84,000.00

$       100,000.00

C = (A/B) x 100

Profit Margin

13.83%

17.00%

Working

Yanik Company

Nunez Compnay

A

Gross Profit

$         26,000.00

$         40,000.00

B

Net Sales

$         84,000.00

$       100,000.00

C = (A/B) x 100

Gross Profit Rate

30.95%

40.00%

  • Requirement ‘c’

--Nunez Company has more sales than Yanik Company

--From ‘b’ analysis, its apparent that Nunez Company is doing better than Yanik Company.

--Nunez’ Sales are more, Gross Profit rate is higher and also the Profit margin is higher than that of Yanik’s.